# Retrieval-Augmented Decision Transformer: External Memory for In-context RL

### v1

@@ 1 Introduction
  • We release datasets for Dark-Room, Dark Key-Door, Maze-Runner, and Procgen to foster future research on in-context decision-making that leverages offline pre-training [^1] GitHub: https://github.com/ml-jku/RA-DT.
@@ 3.2.1 Vector Index for Retrieval Augmentation
In RAG applications for Natural Language Processing (NLP), a common choice for $g(\cdot)$ is a pre-trained LM. While pre-trained models in NLP are ubiquitous, they are rarely available in RL. A natural choice to instantiate $g(\cdot)$ is to train a DT on the pre-collected dataset $\mathcal{D}$, as they exhibit a well-separated embedding space after pre-training [Schmied et al., 2023]. Therefore, they are well suited for retrieval since a new task can be matched to similar tasks in the vector index. As a domain-agnostic alternative, we propose to utilize the FrozenHopfield (FH) mechanism Paischer et al. [2022] to map trajectories to the embedding space of a pre-trained LM. This enables instantiating $g(\cdot)$ with a pre-trained language encoder. The FH mechanism is parameterized by an embedding matrix $\bm{E}\in\mathbb{R}^{v\times d_{\text{LM}}}$ of a pretrained LM with vocabulary size $v$ and hidden dimension $d_{\text{LM}}$, a random matrix $\bm{P}$ with entries sampled from $\mathcal{N}(0,d_{\text{in}}/d_{\text{LM}})$, and a scaling factor $\beta$ and performs:
@@ 4.6 Ablations
Sensitivity of Reweighting. In addition, we conduct a sensitivity analysis on $\alpha$ used in the reweighting mechanism (see Equation 4) that determines the influence of utility on the retrieval score. In Figure 6b, we find that RA-DT performs well for a range of values for $\alpha$ used during training, but performance declines if no re-weighting is employed ($\alpha=0$). We perform the same analysis for $\alpha$ during evaluation in Figure 29.
@@
For additional ablations on RA-DT (query construction, placement of cross-attention layers, retrieval steps) and on our baselines (effect of $K$ in AD), we refer to Appendix E.
@@ 5 Discussion
Memory-Exploitation vs. Meta-learning Abilities. Current offline in-context RL methods are predominantly evaluated on contextual bandits or grid-worlds, such as Dark-Room [Laskin et al., 2022; Lee et al., 2023; Lin et al., 2023; Sinii et al., 2023; Huang et al., 2024], which can only be solved by leveraging the context. However, it remains unclear to what extent the agent learns to learn in-context or simply copies from its context. Further, in our experiments on fully-observable environments (MetaWorld, DMControl, and Procgen), we did not observe ICL behaviour (see Appendices D.3, D.4, D.5). Therefore, it is necessary that future research on in-context RL disentangles the effects of memory and meta-learning abilities, similar to memory and credit-assignment [Ni et al., 2024]. We believe our datasets facilitate future work in this direction.
Challenges of Next-Action Prediction. Most in-context RL methods learn from offline datasets via next-action prediction and causal sequence modelling objectives. As such, they cannot learn to infer the utility of an action, and thus, distinguish between positive and negative examples. This can induce delusions, which lead to repetitions of suboptimal actions and copying behaviour [Ortega et al., 2021] (see Figure 19 for examples on Dark-Room). In contrast, online in-context RL methods have shown promising adaptation abilities [Team et al., 2023; Grigsby et al., 2023; Lu et al., 2024]. Consequently, a potential remedy to this problem is to train a value function to learn the utility of an action, as is commonly done in offline RL [Levine et al., 2020].
@@
Diversity of the Pre-training Distribution. Finally, the diversity and scale of the pre-training dataset may significantly affect the emergence of in-context learning. In our experiments, we pre-train on a relatively small set of tasks. Our results on gridworlds suggest that this is sufficient for ICL to emerge on simple environments. However, on more complex environments, the unseen tasks can be considered out-of-distribution. Therefore, higher pre-training diversity may be necessary for ICL to emerge. It remains unclear how much diversity is required in the pre-training distribution to elicit in-context RL, and if existing large-scale agents exhibit ICL [Reed et al., 2022; Raad et al., 2024]. One promising approach is to expand the pre-training diversity through learned interactive simulations [Yang et al., 2023a; Bruce et al., 2024].
@@ 6 Conclusion
ICL is a promising avenue towards more general agents. Existing in-context RL methods keep entire episodes in their context window, which is challenging as RL environments are typically characterized by long episodes and sparse rewards. To address this challenge, we introduce RA-DT, which employs an external memory mechanism to store past experiences. This enables the retrieval of sub-trajectories that are relevant for the current situation. Our RA-DT outperforms baselines on grid-worlds, while using only a fraction of their context length. Furthermore, we found that RA-DT improves average performance on holdout tasks on complex environments (e.g., robotics and video games). However, RA-DT along with other established in-context RL methods struggles to exhibit in-context improvement. Consequently, we illuminate the current limitations of in-context RL methods and discuss future directions. Finally, we release our datasets for Dark-Room, Dark Key-Door, MazeRunner, and Procgen, to facilitate future research on in-context RL.
Future Work. Besides the general directions discussed in Section 5, we highlight a number of concrete approaches to extend RA-DT. While we focus on in-context improvement without relying on expert demonstrations, pre-filling the external memory with demonstrations may enable RA-DT to perform more complex tasks. This may be particularly powerful for robotics applications, where expert demonstrations are easy to obtain. Furthermore, end-to-end training of the retrieval component in RA-DT, similar to [Izacard et al., 2022], may result in more precise context retrieval and enhanced down-stream performance. Finally, we envision that modern recurrent architectures [Bulatov et al., 2022; Gu & Dao, 2023; Beck et al., 2024] as policy backbones may benefit RA-DT by maintaining hidden states across many episodes.
@@ B.4 DMControl
DMControl contains 30 different robotic tasks with different robot morphologies [Tassa et al., 2018]. Similar to prior work [Hafner et al., 2019; Schmied et al., 2023], we select 16 of these 30 tasks and split them into 11 training (DMC11) and 5 evaluation tasks (DMC5). The DMC11 training tasks are:
@@ E.6 Interaction steps between context retrieval
E.7 Pre-trained Language Model
@@
E.8 Effect of $K$ on Algorithm Distillation

### v2

@@ 1 Introduction
  • We release datasets for Dark-Room, Dark Key-Door, Maze-Runner, and Procgen to foster future research on in-context decision-making that leverages offline pre-training. [^1] GitHub: https://github.com/ml-jku/RA-DT.
@@ 3.2.1 Vector Index for Retrieval Augmentation
In RAG applications for Natural Language Processing (NLP), a common choice for $g(\cdot)$ is a pre-trained LM. While pre-trained models in NLP are ubiquitous, they are rarely available in RL. A natural choice to instantiate $g(\cdot)$ is to train a DT on the pre-collected dataset $\mathcal{D}$, as they exhibit a well-separated embedding space after pre-training [Schmied et al., 2024]. Therefore, they are well suited for retrieval since a new task can be matched to similar tasks in the vector index. As a domain-agnostic alternative, we propose to utilize the FrozenHopfield (FH) mechanism Paischer et al. [2022] to map trajectories to the embedding space of a pre-trained LM. This enables instantiating $g(\cdot)$ with a pre-trained language encoder. The FH mechanism is parameterized by an embedding matrix $\bm{E}\in\mathbb{R}^{v\times d_{\text{LM}}}$ of a pretrained LM with vocabulary size $v$ and hidden dimension $d_{\text{LM}}$, a random matrix $\bm{P}$ with entries sampled from $\mathcal{N}(0,d_{\text{in}}/d_{\text{LM}})$, and a scaling factor $\beta$ and performs:
@@ 4.6 Ablations
Sensitivity of Reweighting. We conduct a sensitivity analysis on $\alpha$ used in the reweighting mechanism (see Equation 4). In Figure 6b, we find that RA-DT performs well for a range of values for $\alpha$ used during training, but performance declines if no re-weighting is employed ($\alpha=0$). We perform the same analysis for $\alpha$ during evaluation in Figure 29.
@@
Effect of Retrieval on Training/Inference Efficiency. Retrieval-augmentation adds computational overhead to the training/inference pipeline due to the cost of embedding the query and searching for similar experiences. However, we find that RA-DT results in significantly faster training times because of shorter context length (up to 7×7$\times$7 × see Appendix E.7). At inference-time, RA-DT is slightly slower compared to baselines when retrieving at every step, but exhibits similar inference speeds when retrieving less frequently (see Appendix E.8). Importantly, the retrieval mechanism in RA-DT enables access to the entirety of the experiences collected across all ICL trials with small additional cost. The ability to access a broader range of experiences may be a reason for its enhanced performance.
For additional ablations on RA-DT and on our baselines, we refer to Appendix E.
@@ 5 Discussion
Memory-Exploitation vs. Meta-learning Abilities. Current offline in-context RL methods are predominantly evaluated on contextual bandits or grid-worlds, such as Dark-Room [Laskin et al., 2022; Lee et al., 2023; Lin et al., 2023; Zisman et al., 2023; Sinii et al., 2023; Huang et al., 2024], which can only be solved by leveraging the context. However, it remains unclear to what extent the agent learns to learn in-context or simply copies from its context. Further, in our experiments on fully-observable environments (MetaWorld, DMControl, and Procgen), we did not observe ICL behaviour (see Appendices D.3, D.4, D.5). Therefore, it is necessary that future research on in-context RL disentangles the effects of memory and meta-learning abilities, similar to memory and credit-assignment [Ni et al., 2024], potentially on recent benchmarks [Nikulin et al., 2023; 2024].
Challenges of Next-Action Prediction. Most in-context RL methods learn from offline datasets via next-action prediction and causal sequence modelling objectives. As such, they cannot learn to infer the utility of an action, and thus, distinguish between positive and negative examples. This can induce delusions, which lead to repetitions of suboptimal actions and copying behaviour [Ortega et al., 2021] (see Figure 19 for examples on Dark-Room). In contrast, online in-context RL methods have shown promising adaptation abilities [Team et al., 2023; Grigsby et al., 2023; Lu et al., 2024]. A similar trend has been observed in online meta-RL methods [Melo, 2022; Shala et al., 2024]. Consequently, a potential remedy to this problem is to train a value function to learn the utility of an action. However, this is usually not straightforward and requires constrained optimization objectives [Zanette et al., 2021; Kumar et al., 2020]. Therefore we leave this approach to future work.
@@
Diversity of the Pre-training Distribution. The diversity and scale of the pre-training dataset may significantly affect the emergence of ICL. In our experiments, we pre-train on a relatively small set of tasks. Our results on gridworlds suggest that this is sufficient for ICL to emerge on simple environments. However, on more complex environments, the unseen tasks can be considered out-of-distribution and higher pre-training diversity may be necessary for ICL to emerge. It remains unclear how much diversity is required to elicit in-context RL, and if existing large-scale agents exhibit ICL [Reed et al., 2022; Raad et al., 2024]. One promising approach is to expand the pre-training diversity through learned interactive simulations [Yang et al., 2023a; Bruce et al., 2024].
@@ 6 Conclusion
Existing in-context RL methods keep entire episodes in their context window, which is challenging as RL environments are typically characterized by long episodes and sparse rewards. To address this challenge, we introduce RA-DT, which employs an external memory mechanism to store past experiences and to retrieve experiences relevant for the current situation. RA-DT outperforms baselines on grid-worlds, while using only a fraction of their context length. While RA-DT improves average performance on holdout tasks on complex environments, it struggles to exhibit ICL, along with other in-context RL methods. Consequently, we illuminate the current limitations of in-context RL methods and discuss future directions. Finally, we release our datasets for Dark-Room, Dark Key-Door, MazeRunner, and Procgen, to facilitate future research on in-context RL.
Future Work. Besides the general directions discussed in Section 5, we highlight a number of concrete approaches to extend RA-DT. While we focus on ICL without relying on expert demonstrations, pre-filling the external memory with demonstrations may enable RA-DT to perform more complex tasks. This may be effective for robotics applications, where expert demonstrations are easy to obtain. Furthermore, end-to-end training of the retrieval component in RA-DT, similar to [Izacard et al., 2022], may result in more precise context retrieval and enhanced down-stream performance. Finally, we envision that modern recurrent architectures [Bulatov et al., 2022; Gu & Dao, 2023; Beck et al., 2024] as policy backbones may benefit RA-DT by maintaining hidden states across many episodes.
@@ B.4 DMControl
DMControl contains 30 different robotic tasks with different robot morphologies [Tassa et al., 2018]. Similar to prior work [Hafner et al., 2019; Schmied et al., 2024], we select 16 of these 30 tasks and split them into 11 training (DMC11) and 5 evaluation tasks (DMC5). The DMC11 training tasks are:
@@ E.6 Interaction steps between context retrieval
E.7 Effect of retrieval-augmentation on Training efficiency
Retrieval-augmentation adds computational overhead to the training pipeline due to the cost of embedding the query trajectories and searching for similar experiences in the vector index. Therefore, we study the effect of retrieval-augmentation on the training efficiency of RA-DT. For the purpose of this analysis, we measure training efficiency in terms of the number of samples processed per second (higher is better). We run all experiments on an A100 GPU using the same training setup (batch sizes, context lengths) as described in Appendix C.
In Figure 32, we compare domain-specific/agnostic RA-DT to the three considered baselines on Dark-Room across gridsizes $10\times 10$, $20\times 20$, and $40\times 20$. We find that the domain-specific variant of RA-DT attains minor training speed-ups on $10\times 10$ and trains almost 7$\times$ faster than baselines on the largest grid. The domain-agnostic variant of RA-DT, in contrast, exhibits slower training times on $10\times 10$, but also trains significantly faster on the largest grid. Note that the differences among the three grid-sizes in the number of samples processed per second of RA-DT stem from the difference in sequence lengths ($C=50$ for $10\times 10$, $C=100$ for $20\times 20$/$40\times 20$) and batch sizes ($B=128$ for $10\times 10$/$20\times 20$, $B=32$ for $40\times 20$).
The efficiency gains of RA-DT are a direct result of the shorter required sequence lengths. In contrast to the baselines, the computational requirements of RA-DT do not grow with the episode length of the environment. Additional speed-ups can be achieved for RA-DT by pre-computing the retrieved trajectories prior to training similar to Borgeaud et al. [2022]. We also want to highlight that all baselines use FlashAttention to speed-up the training times and to ensure a fair comparison. Consequently, the empirical evidence demonstrates that RA-DT does not only improve the down-stream performance in the environments, but is is also significantly faster to train (up to 7×7$\times$7 ×).
(a) Dark-Room 10$\times$10
(b) Dark-Room 20$\times$20
(c) Dark-Room 40$\times$20
Figure 32: Training efficiency for all considered methods on (a) Dark-Room 10$\times$10, (b) Dark-Room 20$\times$20, (c) Dark-Room 40$\times$20. We measure training efficiency in terms of the number of samples processed per second (higher is better). RA-DT achieves considerably speed-ups, in particular for larger grid-sizes.
E.8 Effect of retrieval-augmentation on Inference efficiency
Retrieval-augmentation also incurs computational overhead during inference. Therefore, we study the effect of retrieval-augmentation on the infernce efficiency of RA-DT, similar to Appendix E.7. We measure inference efficiency in the number of environment interaction steps performed per second (higher is better). Note that this metric includes the environment latency. We average the inference efficiency metric across episodes to get a more robust estimate and discard the first episode to exclude compilation times. We conduct our analysis on an A100 GPU and use the same inference setup as described in Appendix C.
In Figure 33, we report the inference efficiency for domain-specific/agnostic RA-DT and the considered baselines on Dark-Room. For RA-DT, we report the inference times with $t\in\{1,25\}$ where $t$ represents the number of interaction steps between retrievals. In Appendix E.6 we found that increasing $t$ only results in minor performance drops for RA-DT. For $t=1$ RA-DT exhibits slightly slower inference speeds compared to the baselines. In contrast, for $t=25$ there is no significant difference in inference speed between RA-DT and the baselines.
(a) Dark-Room 10$\times$10
(b) Dark-Room 20$\times$20
(c) Dark-Room 40$\times$20
Figure 33: Inference efficiency for all considered methods on (a) Dark-Room 10$\times$10, (b) Dark-Room 20$\times$20, (c) Dark-Room 40$\times$20. We measure inference efficiency in terms of the number of environment interaction steps performend per second (higher is better).
Note that the inference speed is roughly the same across grid sizes. This suggests that the inference time is not yet dominated by the quadratic cost of self-attention for $B=1$ and the sequence lengths we consider in this analysis. To further support this, we run an ablation in which we compare the inference efficiency for all baselines with and without FlashAttention (see Figure 35) on Dark-Room $40\times 20$. Indeed, we observe a significant drop in inference speed when FlashAttention is disabled.
Figure 34: Effect of FlashAttention on inference efficiency of AD, DPT and DT on Dark-Room $40\times 20$. Disabling FlashAttention results in a considerable drop in inference speed.
To conclude this analysis, our findings indicate that while RA-DT is slightly slower when retrieving on every step, it achieves comparable inference speeds to the baselines when retrieving less frequently. Importantly, the retrieval mechanism in RA-DT enables access to the entirety of the experiences collected across all ICL trials. In contrast, the baselines can only access experiences from a limit set of the most recent episodes that are preserved in the context (2 in our experiments). If we were to provide more context episodes to the baselines, the quadratic complexity of self-attention would kick in (similar to Figure 32). The ability of RA-DT to access a much broader set of experiences may be a reason for its enhanced down-stream performance.
Figure 35: Effect of FlashAttention on inference efficiency of AD, DPT and DT on Dark-Room $40\times 20$. Disabling FlashAttention results in a considerable drop in inference speed.
E.9 Pre-trained Language Model
@@
E.10 Effect of $K$ on Algorithm Distillation
@@
E.11 Convergence of Baselines
In the main experiments on grid-worlds reported in Section 4.1, we found that the baselines AD and DPT only reach sub-optimal performance within the 40 ICL trials. Therefore, we analyse their performance if evaluate for more ICL trials. In Figure 38, we compare the evaluation performance of AD and DPT across 200 ICL trials on the 20 hold-out tasks for Dark-Room $10\times 10$. We find that both method continue to improve towards optimal performance in this environment when given more ICL trials. For this ablation, we found it useful to set $K=50$ in AD (see Appendix E.10) instead of $K=100$ as used in our main experiments over 40 ICL trials.
Figure 38: Evaluation of AD and DPT on Dark-Room 10$\times$10 over 200 ICL trials. Both methods continue to improve towards optimal performance on this environment.